Find the solution to the system of equations.
x-3y-3z=-4
x+4y+3z=4
y+5z=26

Answer by TimothyLamb(4379)

---
x-3y-3z=-4
x+4y+3z=4
y+5z=26
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x= 2
y= -4
z= 6
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php